How much more output will the average American (U.S population = 330 million) make a year from now if the population does not change but the $20 trillion GDP grows by

a. 0.8 percent

b. 1.5 percent

c. 3 percent

All answers in whole numbers
Answer #1

Case a) When the growth rate is 0.8%

  • Current GDP = 20 trillion
  • Expected GDP = 20 x (1+0.8%)
    • Expected GDP = 20.16 trillion
  • Population = 330 million
  • So the per capital output increase = (20.16-20) trillion/330 million
    • 1 trillion = 1012
    • 1 million = 106
    • So the per capital output increase = $485

Case b) When the growth rate is 1.5%

  • Current GDP = 20 trillion
  • Expected GDP = 20 x (1+1.5%)
    • Expected GDP = 20.30 trillion
  • Population = 330 million
  • So the per capital output increase = (20.30-20) trillion/330 million
    • 1 trillion = 1012
    • 1 million = 106
    • So the per capital output increase = $909

Case c) When the growth rate is 3%

  • Current GDP = 20 trillion
  • Expected GDP = 20 x (1+1.5%)
    • Expected GDP = 20.60 trillion
  • Population = 330 million
  • So the per capital output increase = (20.60-20) trillion/330 million
    • 1 trillion = 1012
    • 1 million = 106
    • So the per capital output increase = $1818